March 20 2009

It was another gorgeous day for measurements today. The day started out calm, clear and very warm up at the lab (-18C!) though the wind picked up and the temperatures dropped by later in the day. PARIS had a few troubles getting started today - we're not sure if it was too hot (the lab was up to 30C) or the vibrations of the Bruker pump upset it (though it doesn't usually), or if it was just random communication errors. With Felicia on the phone though, we were able to go through a series of attempts, and eventually it decided to work normally and measure for the rest of the day. I also installed a red glass filter in the Bruker today, to limit the light if we want to make near infrared (NIR) measurements. After accumulating 70 of or so of our normal mid infrared measurements, I did just that, picking up a handful of NIR measurements late in the afternoon. We are hoping to be able to make NIR measurements in support of the Total Carbon Column Observing Network for validation of GOSAT (a Japanese carbon-dioxide measuring satellite) in the future, but to meet those stringent requirements, we need a new detector which is more sensitive in these wavelengths. The measurements I made with our current detector today do, however, let me experiment with the analysis software needed for those CO2 retrievals.

The UV-VIS instruments continue to work well, though SAOZ was recording some odd-looking spectra in the afternoon. These are back to normal now (Saturday morning) so I'm not sure quite what it was detecting. Yan also implemented a new data transfer procedure for the GBS instruments today, so that data is again being transferred south routinely with the CANDAC data suite.

The station continues to be Grand Central Station, with Defense personnel, UNCLOS people, Kenn Borek Air pilots and engineers and North Pole explorers passing through on a daily or weekly basis. With so many people on station, the fresh food supplies are almost gone - but somehow Jon and Donna manage to keep on feeding us wonderful meals!

Bec




Back to Daily Reports
Back to Home